# A Practical Probabilistic Benchmark for AI Weather Models

### v1

@@ Abstract
Since the weather is chaotic, forecasts aim to predict the distribution of future states rather than make a single prediction. Recently, multiple data driven weather models have emerged claiming breakthroughs in skill. However, these have mostly been benchmarked using deterministic skill scores, and little is known about their probabilistic skill. Unfortunately, it is hard to fairly compare AI weather models in a probabilistic sense, since variations in choice of ensemble initialization, definition of state, and noise injection methodology become confounding. Moreover, even obtaining ensemble forecast baselines is a substantial engineering challenge given the data volumes involved. We sidestep both problems by applying a decades-old idea—lagged ensembles—whereby an ensemble can be constructed from a moderately-sized library of deterministic forecasts. This allows the first parameter-free intercomparison of leading AI weather models’ probabilistic skill against an operational baseline. The results reveal that two leading AI weather models, i.e. GraphCast and Pangu, are tied on the probabilistic CRPS metric even though the former outperforms the latter in deterministic scoring. We also reveal how multiple time-step loss functions, which many data-driven weather models have employed, are counter-productive: they improve deterministic metrics at the cost of increased dissipation, deteriorating probabilistic skill. This is confirmed through ablations applied to a spherical Fourier Neural Operator (SFNO) approach to AI weather forecasting. Separate SFNO ablations modulating effective resolution reveal it has a useful effect on ensemble dispersion relevant to achieving good ensemble calibration. We hope these and forthcoming insights from lagged ensembles can help guide the development of AI weather forecasts and have thus shared the diagnostic code [^1] http://github.com/NVIDIA/earth2mip.
@@ 1 Introduction
In recent years, several data-driven weather models have claimed deterministic skill that are comparable (and even outperform) state of the art numerical weather models. AI weather models are very computationally efficient, suggesting that they could be ideal for probabilistic forecasting with large ensembles.
@@
Unfortunately, little is reported about the probabilistic skill of recent AI weather models. In contrast, deterministic skill scores have been widely reported – indeed, most AI forecasting systems have been optimized using either mean squared error (Bonev et al., 2023; Lam et al., 2023) or mean absolute error (Bi et al., 2023; Chen et al., 2023b) as a key part of the loss function, and correspondingly their performance has often been compared to state-of-the-art deterministic forecasts (Bi et al., 2023; Bonev et al., 2023; Chen et al., 2023b; Lam et al., 2023). Yet this is unsatisfying as the Root Mean Squared Error (RMSE) for a forecast from a perfect deterministic model saturates at a value $\sqrt{2}$ higher than that for a perfect ensemble forecast as derived theoretically by Leith (1974) and demonstrated for an ensemble of AI weather models by Weyn et al. (2021). Any model claiming better deterministic skill at long lead times does so by artificially reducing the variance of the forecast, which will lead it to underestimate the chances of unlikely events.
@@
The purpose of this paper is to present a consistent and easy-to-use method for evaluating deterministic weather forecasts from the lens of probabilistic skill; to show how this clarifies the pitfalls of using only deterministic scoring for model comparison, and to add complementary perspective to ongoing discussions about how best to optimize AI weather predictions. We speculate the lack of such a tool underlies the dearth of probabilistic score intercomparisons in the literature, and stems from three key challenges.
The first is that fair comparison against an operational ensemble is an overwhelming technical challenge, at least if one considers the typical requirements of this baseline. The common operational practice is to run an ensemble of deterministic forecasts to estimate the probability distribution of future weather states. In such ensemble forecasts, the initial state (and sometimes the model) are perturbed to account for unavoidable uncertainties. Model-aware initialization techniques tend to require some knowledge of the forecast model to come up with initial conditions at a time $t$, potentially from the past history. For instance, some ensemble techniques like vector breeding Toth and Kalnay (1997) require running the model while others like singular value decomposition require knowledge of the model’s adjoint (Molteni et al., 1996; Leutbecher, 2005). Model-dependent noise-injection techniques are further used to enhance dispersion but to varying degrees across operational centers Palmer et al. (2009). Perturbation techniques for traditional weather models cannot be straightforwardly applied to AI weather models. In short, fair comparison of deterministic weather predictions against such an ensemble forecasting system is difficult as the perturbation techniques and their tunable parameters are often model specific. The ensemble methods used in operational centers have been tuned continuously over decades, and there is no uniform ensemble perturbation method across numerical weather prediction centers. That is, a parameter free intercomparison method would be attractive.
@@
The second challenge is that scoring ensemble forecasts over many lead times requires large data volumes that are not easily transferred between various groups that develop models. The amount of data of a modern ensemble of 50-100 members is prohibitive. Assuming we verify at $0.25^{\circ}$ resolution, with 12-hour sampling up to 14 days, and 600 initial times, then the baseline data needed for scoring for a single ensemble member of one scalar field (e.g. geopotential height at 500 hPa) is 70GB. For 50 members, this becomes TBs, whereas the operational archives often have a peak download speed of around 10 MB/s. Thus, especially with the increasing demand from the growing community experimenting with AI weather prediction, simply attaining the operational baseline is a challenge. This data volume barrier will only grow higher as the resolution of global forecasts increases.
@@
In summary, a solution having the following requirements would enable more systematic probabilistic scoring of forecasts (either data-driven or traditional).
  1. 1.
    Use the same ensembling method for all models. Remove confounding factors.
  2. 2.
    Minimize data requirements.
  3. 3.
    Minimize tuning knobs. Keep it simple.
  4. 4.
    Allow rapid feedback to the AI model developer.
We suggest such a solution already exists: A decades-old idea - lagged ensemble forecasts (LEF) - can be used as a strategic ensemble of convenience.
@@
Practically, LEF have several properties that make them ideal for validating and intercomparing the deterministic forecast models that form the core of an operational ensemble forecast system. LEF provides a parameter free ensemble forecasting system (Branković et al., 1990). As such it can be used to fairly compare between different numerical and machine learning weather models. Lagged ensembles allow one to compute a probabilistic score from deterministic forecast archives, which are not prohibitive in size and can reasonably be moved around the internet or computed on the fly from data-driven models. The method is extremely simple, and most importantly the same for all models we consider, though there may be some differences in the analyses used to start the deterministic forecasts. The technique is especially valuable for comparing ML weather models consistently against physical model baselines.
In the following sections, we apply LEF to intercompare the IFS and leading AI weather models for the first time, to draw several insights. First, we validate the relevance of LEF by showing that, applied to IFS, it can explain most of the temporal variance of the actual operational ensemble (Sec. 3.1). Then we apply LEF to re-examine the relative skill of two leading AI weather models, to illustrate that a particular AI intervention – multi-step fine-tuning - that leads to apparent gains in deterministic skill yields no associated gains in probabilistic skill, consistent with the speculation of (Lam et al., 2023) (Sec. 3.2). We then confirm this hypothesis by testing the effects of multi-step fine-tuning across a detailed set of ablations with the spherical Fourier neural operator (SFNO) AI weather model Bonev et al. (2023) (Sec. 3.3). We then turn to assessing the role of effective resolution in the context of SFNO architectures where we find that, similarly to fine tuning, systematically decreasing the effective resolution also reduces the spread of the lagged ensemble. Together, these results support our overall thesis – that LEF provides efficient, practical guidance on probabilistic skill, with potential to accelerate the development of state of the art probabilistic weather forecasting systems.
@@ 2 Lagged Ensemble Benchmark
Figure 1: Overview of lagged ensemble forecasting. Each ensemble member (color) is initialized at a different initial time (dots). The true time series (-) and the lagged ensemble average (- -) are also shown.
Figure 1 shows a schematic of the lagged ensemble approach on a simple chaotic dynamical system (see Sec A.1 for more information). The right-hand side shows how forecasts initialized from different initial times soon diverge from one another, when evaluated at common validation time. The ensemble mean has a lower variance, but is clearly too smooth and unlike the truth.
@@
$x_{m}(t,t^{\prime})=x(t,t^{\prime}-mh),\quad-M\leq m\leq M.$
This ensemble has $12M+$ 1 members and the initial times are centered around $t^{\prime}$. While this ensemble cannot be used for a real weather forecast since it requires initial conditions from the future, it is suitable for comparing hindcasts. Assuming $t$ and $t^{\prime}$ are uniformly sampled with $h$ simplifies the implementation. In practice a lagged ensemble can be generated from any archive of deterministic hindcasts (Rasp et al., 2023a; Swinbank et al., 2016), allowing probabilistic scoring on data volumes far smaller than raw ensemble outputs, which Price et al. (2023) pointed out are prohibitive to obtain.
@@
The lagged ensemble (and deterministic) forecasts are compared with the corresponding atmospheric states given by ERA5 reanalysis Hersbach et al. (2020) at $0.25^{\circ}$ resolution, which is also the training data for all the ML models evaluated in this study. Hindcasts are made with initial times taken every 12 hours from 2018-01-02 0Z through 2018-10-29 12Z, which is outside the training set for all the models we consider. Like IFS HRES, we consider 10-day forecasts.
There are some important caveats about our choice of ERA5 as verification. First, to fairly evaluate models with different spatial resolution, we should down-sample to the coarsest available resolution (Rasp et al., 2023b). Also, ERA5 dataset provides fictitious numerical values on pressure levels that lie below the surface, based on extrapolating a field downwards from the surface in a field-dependent way. It does not seem fair to evaluate models on how well they predict these unphysical values. We sidestep these concerns in the present work by only considering models trained with these same data so we leave it to future iterations to addresses these concerns.
@@
We use standard metrics to evaluate the performance of a given lagged ensemble. First, a summary metric of the probability distribution penalizing both over-confidence and inaccuracy is measured using the continuous-ranked-probability score (CRPS)(Gneiting and Raftery, 2007). Second, we evaluate the accuracy of the ensemble mean using the root-mean-square-error (eRMSE, i.e. ensemble-RMSE). Third, we define the deterministic RMSE (dRMSE) as the RMSE of the $m=0$ ensemble member. Finally, we diagnose how well calibrated a lagged ensemble is by comparing the growth of the ensemble standard deviation (spread) to the growth of the eRMSE. One ensemble is more dispersive than another if its spread is higher for a given level of error (eRMSE). All metrics are globally averaged and area-weighted. Section A.4 contains more details.
For simplicity we assume all ensemble members are equally weighted. For short lead time forecast, assigning larger weights to more recent lags is preferable as in (Hoffman and Kalnay, 1983), but for the purpose of having a parameter-free benchmark, we propose using uniform weights across the ensemble. This somewhat complicates the spread-error and other calibration analyses since the famous 1:1 spread-error relationship assumes ensemble members are interchangeable with each other and the observations (Fortin et al., 2014). Nonetheless, we do use the spread-error relationship to diagnose if an ensemble is more or less dispersive than another, but will take care to avoid value judgements like “under” and “over” dispersive. We will sometimes use the IFS results as a target since the ECWMF operational forecast ensemble built on top of IFS is well-calibrated. Section A.5 elaborates how we could extend this analysis to weighted ensembles.
@@ 3.1 Comparing a lagged and operational ensemble
It is natural to wonder to what extent the LEF method can be used as a valid proxy of the operational ensemble. To validate this idea, we compare the lagged ensemble of IFS against the ECMWF operational ensemble ECMWF (2019a), see also A.2. By comparing the best deterministic and ensemble forecasting systems from within a single modeling center, we remove many confounding factors in the analysis, although it is worth noting that the IFS and HRES ensembles were generated at different resolutions and with different configurations.
Figure 2 confirms the validity of LEF as a proxy with meaningful operational relevance. At a lead time of 5 days, i.e. sufficiently long that no member of the LEF contains the observations, the CRPS of the lagged and operational ensembles are well-correlated across initial dates: much of the day-to-day variation in global CRPS of the operational ensemble can be predicted from that of the lagged ensemble. As expected, the lagged ensemble is more skillful on average, since it has the artificial advantage that some of its members have shorter lead times than the operational ensemble. We have confirmed that the relationship between LEF CRPS and operational CRPS is similarly monotonic for other variables and shorter lead times (not shown).
In short, even though there are many differences between these ensemble methods, Figure 2 supports our working hypothesis that LEF is a meaningful probabilistic benchmark of operational relevance. We thus proceed to employ it to perform the first clean probabilistic intercomparison of AI weather models against the IFS LEF baseline.
[…]

### v2

@@ Abstract
Since the weather is chaotic, it is necessary to forecast an ensemble of future states. Recently, multiple AI weather models have emerged claiming breakthroughs in deterministic skill. Unfortunately, it is hard to fairly compare ensembles of AI forecasts because variations in ensembling methodology become confounding and the baseline data volume is immense. We address this by scoring lagged initial condition ensembles—whereby an ensemble can be constructed from a library of deterministic hindcasts. This allows the first parameter-free intercomparison of leading AI weather models’ probabilistic skill against an operational baseline. Lagged ensembles of the two leading AI weather models, GraphCast and Pangu, perform similarly even though the former outperforms the latter in deterministic scoring. These results are elaborated upon by sensitivity tests showing that commonly used multiple time-step loss functions damage ensemble calibration.
Plain Language Summary
2023 was a seminal year for data-driven weather forecasts with several high-profile publications claiming that AI outperformed traditional physics-based approaches to weather forecasts. These claims are mostly supported by scoring deterministic forecasts, even though it is widely known that forecasting is a probabilistic problem. Probabilistic intercomparisons have proved challenging because of the data volumes involved and because they are confounded by particulars of how probabilistic forecasts are built. As a workaround, we propose benchmarking weather forecasts using lagged ensemble forecasting where forecasts initialized at different times are considered independent samples. When benchmarked in this way, we show that some AI models achieve better deterministic scores by reducing the variance of their forecasts at the cost of physical realism.
@@ 1 Introduction
Recently data-driven weather models have begun to outperform state-of-the-art physics-based models both in terms of deterministic skill (Lam et al., 2023; Bi et al., 2023; Chen et al., 2023a, b) and probabilistic forecasting with large ensembles (Price et al., 2023; Kochkov et al., 2023; Zhong et al., 2024; Mahesh et al., 2024).
Despite these recent successes in probabilistic forecast, benchmarking probabilistic AI models remains a substantial challenge due to the latency of scoring large hind-cast ensembles and the difficulty obtaining baseline data. On the one hand, large data repositories like WeatherBench 2.0 (Rasp et al., 2023) and TIGGE (Swinbank et al., 2016) have greatly improved access to baseline data though not eliminated the challenge. However the latency issue is serious since ML model development is an inherently iterative process requiring rapid feedback on the end-to-end performance of the models. It is common to perform multiple evaluations a week or even during a single model training. In many ways (e.g. storage, I/O), a full ensemble scoring is a more substantial computation than the training itself, so by the time one model is scored the rest of the team has moved on.
@@
There is also a need for benchmarking deterministic forecasts from the lens of probabilistic skill. Even the intrinsically probabilistic models like Gencast (Price et al., 2023) have been built off strong deterministic backbones like Graphcast (Lam et al., 2023), so we may expect that improving the realism of individual ensemble members could improve the overall probabilistic forecast. However, scores like root-mean-squared-error (RMSE) are easy to improve trivially by reducing the variance of a forecast. By training on mean squared error (Bonev et al., 2023; Lam et al., 2023) or mean absolute error (Bi et al., 2023; Chen et al., 2023b) as a key part of the loss function, and correspondingly their performance has often been compared to state-of-the-art deterministic forecasts (Bi et al., 2023; Bonev et al., 2023; Chen et al., 2023b; Lam et al., 2023). Yet this is unsatisfying as the Root Mean Squared Error (RMSE) for a forecast from a perfect deterministic model saturates at a value $\sqrt{2}$ higher than that for a perfect ensemble forecast as derived theoretically by Leith (1974) and demonstrated for an ensemble of AI weather models by Weyn et al. (2021). There has been substantial confusion in the literature around this topic. For example, Graphcast relied on variance reduction via multistep finetuning to improve its scores relative to IFS, but the approach was discarded in the follow-up paper GenCast. Any model claiming better deterministic skill at long lead times does so by artificially reducing the variance of the forecast, which will lead it to underestimate the chances of unlikely events.
@@
The purpose of this paper is to present a consistent and easy-to-use method for evaluating deterministic weather forecasts from the lens of probabilistic skill; to show how this clarifies the pitfalls of using only deterministic scoring for model comparison, and to add complementary perspective to ongoing discussions about how best to optimize AI weather predictions. We aim to address the three challenges.
The first is the difficulty of attributing changes in performance to backbone deterministic model rather than the surrounding ensembling framework. Our motivation to control the ensembling system is similar to Magnusson et al. (2022), who disentangled the performance of the data assimilation system from model physics by comparing forecasts initialized with the same state. For ML models the deterministic backbone will typically be a neural network trained on a regression task while for physics it would be either a deterministic GCM or its dynamical core. Traditional forecast models benefit from the modularity of physics and idealized benchmarks of individual components such as the dynamical core (Jablonowski and Williamson, 2006), but ML methods can only be tested end-to-end which confounds comparisons given the assortment of frameworks in practice both for representing uncertainty in the initial condition (Toth and Kalnay, 1997; Molteni et al., 1996; Leutbecher, 2005) and the model itself (Palmer et al., 2009). As for ML forecast systems, there are similarly a large number of training techniques such as variational auto-encoders (Zhong et al., 2024), de-noising diffusion (Price et al., 2023), multi checkpoint ensembles ((Mahesh et al., 2024; Weyn et al., 2021)) being developed in tandem with specific ML architectures. Even with the progress in intrinsically stochastic methods, it seems preferable that the deterministic backbone be as realistic as possible. So it is hard to know which architecture is the most parameter efficient or has the best performance-compute trade-off. This is as opposed to other ML domains such as computer vision where both the training technique (e.g loss function) and evaluations are more standardized, and as a result small changes in accuracy can be attributed to the architecture. Therefore, a parameter-free and quantitative scoring technique specifically targeted at the deterministic backbone is attractive.
@@
The second challenge is that scoring ensemble forecasts over many lead times requires large data volumes that are not easily transferred between various groups that develop models. The amount of data of a modern ensemble of 50-100 members is prohibitive. Assuming we verify at $0.25^{\circ}$ resolution, with 12-hour sampling up to 14 days, and 600 initial times, then the baseline data needed for scoring for a single ensemble member of one scalar field (e.g. geopotential height at 500 hPa) is 70GB. For 50 members, this becomes TBs, whereas we experience download speeds of 10 MB/s for TIGGE(Swinbank et al., 2016) and 100 MB/s from WeatherBench 2.0 (Rasp et al., 2023). Thus, especially with the increasing demand from the growing community experimenting with AI weather prediction, simply obtaining the operational baseline is a challenge. This download cost can be somewhat mitigated by evaluating at coarser resolution like 1.5 degrees or by scoring with the identical scoring routines, but both of these have their limitations. This data volume barrier will only grow higher as the resolution of global forecasts increases.
@@
In this article, we propose a simple method for benchmarking that addresses all of this challenges, and is suitable for the day-to-day development of improving deterministic backbone architectures and comparing them with existing archives of forecast data. Specifically, we propose a decades-old idea - lagged ensemble forecasts (LEF) - as convenient ensemble for rapid benchmarking.
@@
Practically, LEF have several properties that make them ideal for validating and intercomparing the deterministic forecast models that form the core of an operational ensemble forecast system. LEF provides a parameter free ensemble forecasting system (Branković et al., 1990). As such it can be used to fairly compare between different numerical and machine learning weather models. Lagged ensembles allow one to compute a probabilistic score from deterministic forecast archives, which are not prohibitive in size and can reasonably be moved around the internet or computed on the fly from data-driven models. The method is simple, and most importantly the same for all models we consider, though there may be some differences in the analyses used to start the deterministic forecasts. The technique is especially valuable for comparing ML weather models consistently against physical model baselines.
In the following sections, we apply LEF to intercompare the IFS and leading AI weather models for the first time, to draw several insights. First, we validate the relevance of LEF by showing that, applied to IFS, it can explain most of the temporal variance of the actual operational ensemble (Sec. 3.1). Then we apply LEF to re-examine the relative skill of two leading AI weather models, to illustrate that a particular AI intervention – multi-step fine-tuning - that leads to apparent gains in deterministic skill yields no associated gains in probabilistic skill, consistent with the speculation of (Lam et al., 2023) (Sec. 3.2). We then confirm this hypothesis by testing the effects of multi-step fine-tuning across a detailed set of ablations with the spherical Fourier neural operator (SFNO) AI weather model (Bonev et al., 2023) (Sec. 3.3). We then turn to assessing the role of effective resolution in the context of SFNO architectures where we find that, similarly to fine tuning, systematically decreasing the effective resolution also reduces the spread of the lagged ensemble. Together, these results support our overall thesis – that LEF provides efficient, practical guidance on probabilistic skill, with potential to accelerate the development of state of the art probabilistic weather forecasting systems.
@@ 2 Lagged Ensemble Benchmark
(a) (b)
Refer to caption Refer to caption
[…]
Figure 1a shows a schematic of the lagged ensemble approach on a simple chaotic dynamical system (see Sec A.1 for more information). The right-hand side shows how forecasts initialized from different initial times soon diverge from one another, when evaluated at common validation time. The ensemble mean has a lower variance, but is clearly too smooth and unlike the truth.
@@
$x_{m}(t,t^{\prime})=x(t,t^{\prime}-mh),\quad-M\leq m\leq M.$ (1)
This ensemble has $12M+$ 1 members and the initial times are centered around $t^{\prime}$. While this ensemble cannot be used for a real weather forecast since it requires initial conditions from the future, it is suitable for comparing hindcasts. Assuming $t$ and $t^{\prime}$ are uniformly sampled with $h$ simplifies the implementation. In practice a lagged ensemble can be generated from any archive of deterministic hindcasts (Rasp et al., 2023; Swinbank et al., 2016), allowing probabilistic scoring on data volumes far smaller than raw ensemble outputs, which Price et al. (2023) pointed out are prohibitive to obtain.
@@
The lagged ensemble (and deterministic) forecasts are compared with the corresponding atmospheric states given by ERA5 reanalysis (Hersbach et al., 2020) at $0.25^{\circ}$ resolution, which is also the training data for all the ML models evaluated in this study. Hindcasts are made with initial times taken every 12 hours from 2018-01-02 0Z through 2018-10-29 12Z, which is outside the training set for all the models we consider. Like IFS HRES, we consider 10-day forecasts.
There are some important caveats about our choice of ERA5 as verification. First, to fairly evaluate models with different spatial resolution, we should down-sample to the coarsest available resolution (Rasp et al., 2023). Also, ERA5 dataset provides fictitious numerical values on pressure levels that lie below the surface, based on extrapolating a field downwards from the surface in a field-dependent way. It does not seem fair to evaluate models on how well they predict these unphysical values. We sidestep these concerns in the present work by only considering models trained with these same data so we leave it to future iterations to addresses these concerns.
We use standard metrics to evaluate the performance of a given lagged ensemble. First, a summary metric of the probability distribution penalizing both over-confidence and inaccuracy is measured using the continuous-ranked-probability score (CRPS)(Gneiting and Raftery, 2007). Second, we evaluate the accuracy of the ensemble mean using the root-mean-square-error (eRMSE, i.e. ensemble-RMSE). Third, we define the deterministic RMSE (dRMSE) as the RMSE of the $m=0$ ensemble member. Finally, we diagnose how well calibrated a lagged ensemble is by comparing the growth of the ensemble standard deviation (spread) to the growth of the eRMSE. One ensemble is more dispersive than another if its spread is higher for a given level of error (eRMSE). All metrics are globally averaged and area-weighted. We use unbiased estimators when applicable for CRPS (Zamo and Naveau, 2018), the ensemble standard deviation (Spread), and spread-error comparisons (Fortin et al., 2014). For ensemble mean RMSE, we follow the common practice and report the typical version of the formula. This is biased in favor of under-dispersive models for small ensembles but no clear alternative is documented in the literature. Section A.4 contains more details.
For simplicity we assume all ensemble members are equally weighted. For short lead time forecast, assigning larger weights to more recent lags is preferable as Hoffman and Kalnay (1983) do, but for the purpose of having a parameter-free benchmark, we propose using uniform weights across the ensemble. This somewhat complicates the spread-error and other calibration analyses since the famous 1:1 spread-error relationship assumes ensemble members are interchangeable with each other and the observations (Fortin et al., 2014). Nonetheless, we do use the spread-error relationship to diagnose if an ensemble is more or less dispersive than another, but will take care to avoid value judgements like “under” and “over” dispersive. We will sometimes use the IFS results as a target since the ECWMF operational forecast ensemble built on top of IFS is well-calibrated. Section A.5 elaborates how we could extend this analysis to weighted ensembles.
@@ 3.1 Comparing a lagged and operational ensemble
It is natural to wonder to what extent the LEF method can be used as a valid proxy of the operational ensemble. To validate this idea, we compare the lagged ensemble of IFS against the ECMWF operational ensemble (ECMWF, 2019a), see also A.2. By comparing the best deterministic and ensemble forecasting systems from within a single modeling center, we remove many confounding factors in the analysis, although it is worth noting that the IFS and HRES ensembles were generated at different resolutions and with different configurations.
Figure 1b confirms the validity of LEF as a proxy with meaningful operational relevance. At a lead time of 5 days, i.e. sufficiently long that no member of the LEF contains the observations, the CRPS of the lagged and operational ensembles are well-correlated across initial dates: much of the day-to-day variation in global CRPS of the operational ensemble can be predicted from that of the lagged ensemble. As expected, the lagged ensemble is more skillful on average, since it has the artificial advantage that some of its members have shorter lead times than the operational ensemble. We have confirmed that the relationship between LEF CRPS and operational CRPS is similarly monotonic for other variables and lagged ensemble sizes $M$ (cf. Figure A1) and shorter lead times (not shown).
In short, even though there are many differences between these ensemble methods this analysis supports our working hypothesis that LEF is a meaningful probabilistic benchmark of operational relevance. We thus proceed to employ it to perform the first clean probabilistic intercomparison of AI weather models against the IFS LEF baseline.